The September 11, 2001, terrorist attack on the World Trade Center (WTC) complex damaged or destroyed 21 buildings, accounting for 31.2 million square feet of Manhattan office space – nearly 10 percent of the city’s entire inventory. Suppose demand for office space in New York remained unchanged despite the incident.

Use the following information about the rent for the Manhattan office to answer the next three questions.

The demand for the office is : Qd = 120 – P

The supply of office is: Qs = 3P

Question 2: What is the equilibrium quantity (i.e. space in million square feet) of office rented ?

a. 25

b. 50

c. 75

d. 90
Answer #1

Option d is correct

At equilibrium we have quantity demanded equal to quantity supplied

120 - p = 3p

Now p = 120/4 = 30

This gives Q = 3*30 = 90

Beautiful equilibrium quantity of office rented is 90 million square feet.